# On the Robustness of Verbal Confidence of LLMs in Adversarial Attacks

### v1

@@ C.1 ConfidenceTriggers Settings
To optimize the performance of ConfidenceTriggers, we evaluated a variety of algorithm parameters. However, given the significant amount of iterations necessary to optimize the algorithm, we limited some parameters like trigger prompt length while still aiming for strong results. Regarding our training text corpus, for MedMCQA and StrategyQA, we use 70% of the data for optimizing the triggers, and evaluate it on the remaining 30% (out of which 200 random samples are selected). Given the relatively smaller size of TruthfulQA, we train the triggers on questions with five multiple choices options (yielding 197 different examples) and evaluate on the 203 questions with four options. For our trigger prompt length $\beta$ we choose 20 tokens, and for the total number of prompts we attempt to optimize $\alpha$ we also choose 20. We experimented with optimizing between 20 and 50 total generations. For the number of samples for the initial estimation $S$, we choose 200, and the value $\xi$ for each iteration we choose 12, leading to 48 different sampled training cases used for estimation for each prompt each generation. We choose a fifth of prompts (i.e., 4) with the lowest loss to be elites. An example of the generated triggers can be found in Appendix U.
Our procedure for Cross-Over and Mutation functions remains largely faithful to Lapid et al. [32] and are standard implementations for these functions given two parent sequences of data. Cross-Over is the standard Single point Cross-Over [23] operation and Mutation is Random (uniform) Mutation [17] where words at randomly chosen indices are replaced. Pseudo code for Cross-Over and Mutation can be found in Algorithm 3 and 4 respectively. The tournament select process is top-K like in Blickle and Thiele [4].
@@
4:$\mathbf{C}_{1}$ = $\mathbf{P}_{1}[:SplitPoint]+\mathbf{P}_{2}[SplitPoint:]$
5:$\mathbf{C}_{2}$ = $\mathbf{P}_{2}[:SplitPoint]+\mathbf{P}_{1}[SplitPoint:]$
@@ Appendix D ConfidenceTriggers-AutoDAN Details
In this section, we present the algorithm along with the details for the AutoDAN variation of the ConfidenceTriggers algorithm. The general algorithm can be found in Algorithm 5 and generally follows the original version of ConfidenceTriggers but is modified to incorporate the steps and methods of the AutoDAN-HGA algorithm in Liu et al. [37], in that we make use of the hierarchal structure of prompts and incorporate both sentence and paragraph level optimizations.
@@
[…]
@@
A total of 100 generations is used to optimize the prompts. The remaining settings are consistent with those detailed in Appendix C.1. An example of the final generated triggers can be found in Appendix U.
@@
5:    $\mathcal{L}_{z}$leftarrow AverageConfidence($z$, $S$) $\triangleright$ Base confidence w/o any trigger
@@
7:$P$leftarrow InitializePrompts($\alpha$)
@@
12:             $\mathcal{L}_{p_{g}}$ $\leftarrow\mathcal{L}_{p_{g}}\mathbin{\|}$ AverageConfidence($p$, $z$, $\xi$) -- $\mathcal{L}_{z}$
@@
31:         $p_{a}$leftarrow Mutation($p_{a}$)
32:         $p_{b}$leftarrow Mutation($p_{b}$)
@@
35:    if $|\mathbf{P}_{g}|$ $>$ $|\mathbf{P}|$ then
@@
38:    $\mathbf{P}$leftarrow $\mathbf{P}_{g}$
@@
47:$\mathbf{T}$leftarrow $\mathbf{P}[argmin(\mathcal{L}_{p_{G}})]$triangleright Prompt with lowest loss
@@ Appendix E Confidence Agnostic Attacks
SC 94.6 90.7 100.0 85.5 64.1
We examine if CEMs are robust against generic confidence-agnostic adversarial attacks. To this end, we compare the confidences generated for the original instances in SST-2 [57] and their AdvGLUE [61] versions since these comprehensive adversarial modifications were solely optimized for misclassifications. The results are in Table 7, where the difference in confidences are minor, with at most a 5% drop despite a high percentage of samples being effected. CEMs are shown to be generally robust against confidence-agnostic perturbations, even though they are effective in reducing accuracy by 15-20%.
@@ Appendix F Full Results Tables
SC 91.6 87.1 64.0 7.1 18.0 64.1 91.1 87.9 44.5 6.1 11.7 37.9
[…]
CoT 91.5 83.8 51.0 15.0 38.6 61.8 91.6 85.9 38.5 14.9 12.2 30.4
MS 80.1 64.9 66.5 22.8 31.8 63.2 79.5 65.7 66.5 20.6 20.5 44.1
SC 91.9 84.4 62.5 12.0 37.0 61.5 91.6 86.0 58.5 9.7 20.5 42.6
Avg. - 6.6 46.7 11.9 25.0 58.4 - 4.8 37.9 11.1 13.3 32.5
@@ Appendix H Confidence Phrases
Overall, we see that model behaviour when prompted through user queries varies greatly, but generally responds to be requests and results in great differences in confidence, particularly with the Base CEM. CoT and SC are in general overconfident compared to all methods regardless of how the model is prompted. The level of affected samples is very high, but is not 100% in the majority of cases, hence even this method of directly prompting the model is not always effective at changing model confidence behaviour. Targeting the system prompt is less effective than targeting the user query, particularly with GPT-3.5. The models are likely in most cases ignoring the prompt contents as it is not core to their task.
Table 10: Effectiveness of different phrases appended to user queries. Averages Δ Overall Cf are shown for cases where confidence goes up and down respectively. Negative values represent increases in average confidence compared to the original examples. Averages for Δ Overall Cf are separated by where the confidence goes down and up respectively.
@@
Phrase Adv. Cf Δ Overall Cf % Aff. Δ Aff. Cf %Flip(OC) %Flips(OW) Adv. Cf Δ Overall Cf % Aff. Δ Aff. Cf %Flip(OC) %Flips(OW)
@@
SC 64.8 21.8 97.5 22.4 69.7 76.6 54.0 39.4 99.5 39.6 3.1 44.4
Avg. - 21.0/-6.7 91.7 14.2 72.1 53.5 - 26.8/-6.5 87.4 18.7 9.3 31.5
Table 11: Effectiveness of different phrases appended to system prompts. Averages Δ Overall Cf are shown for cases where confidence goes up and down respectively. Negative values represent increases in average confidence compared to the original examples. Averages for Δ Overall Cf are separated by cases where the confidence goes down and up respectively.
@@
Adv. Cf Δ Overall Cf % Aff. Δ Aff. Cf %Flip(OC) %Flips(OW) Adv. Cf Δ Overall Cf % Aff. Δ Aff. Cf %Flip(OC) %Flips(OW)
@@
SC 82.2 4.8 91.5 5.3 86.8 74.7 88.9 4.5 82.5 5.4 4.7 16.9
Avg. - 8.8/-3.4 73.0 5.5 65.9 55.4 - 6.2/-1.5 61.3 5.5 6.1 21.9
@@ Appendix I Confidence Expressed Via Likert Scale
In addition to numerical confidence scores, it is possible to prompt the model to produce confidence scores in the form of phrases (e.g., very confident) derived from a Likert scale, which can potentially alter model behaviour and bias compared to numeric scores given that phrases are more naturalistic and resemble how humans typically express confidence in the real-world. We aim to analyze how attack performance changes when a model is asked to provide a confidence based on a seven point Likert scale where the options are the following:
@@
The statements are converted to numerical confidence scores for use in attack algorithms by interpolation. The first option is set to max confidence (100%) and the last to maximum entropy ($1/num\_mc\_options*100\%$), and the remaining values are interpolated between these two values according to a linear scale. We determine how effective each of our attack algorithms are in this context, and Table 12 presents the key results. Overall, there is not a noticeable improvement in attack performance compared to when the LLMs produce numerical scores, and performance is very similar in terms of average confidence drop and percent of samples that undergo a drop. We do see that Δ Aff. Cf is very but this can be attributed due to the discrete nature of the scale meaning any change in answer results in a large confidence drop (i.e., it cannot be 5% like when using numerical verbalized scores). In general, however, we do see a similar level of invariance in terms of confidence here and as such this behaviour of the model generalizes across both numerical and phrase-like confidence estimates which means LLMs model both numerical and phrasal confidence similarly and present rigidness in both.
Cf. Adv. Cf % Aff. Δ Aff. Cf %Flip(OC) %Flips(OW)
Llama-3-8B VCA-TF Base 77.9 73.9 7.5 53.3 1.9 9.8
CoT 87.1 84.4 6.0 45.0 10.5 13.2
MS 31.6 23.6 42.0 19.0 19.5 33.6
SC 84.8 77.8 34.5 20.5 39.4 26.7
VCA-TB Base 77.9 73.6 8.0 54.7 2.8 9.8
CoT 87.1 82.3 12.0 39.8 15.1 14.0
MS 31.6 22.1 48.5 19.6 24.1 35.4
SC 85.2 76.1 40.0 22.8 32.3 31.7
SSR Base 77.9 70.4 17.0 44.6 25.9 39.1
CoT 87.1 78.3 22.5 39.4 22.1 50.9
MS 31.6 21.7 56.0 17.6 29.9 54.0
SC 84.3 73.0 61.0 18.4 28.0 57.9
Ty. Base 77.9 74.9 7.0 43.8 5.6 21.7
CoT 87.1 84.5 5.5 47.9 10.5 20.2
MS 31.6 27.8 27.5 13.7 11.5 25.7
SC 84.0 80.7 22.0 15.3 14.9 26.4
Avg. - 6.2 26.1 32.2 18.4 29.4
GPT-3.5 VCA-TF Base 84.5 80.6 12.5 31.3 3.1 8.5
CoT 85.9 85.5 2.0 21.5 0.0 5.3
MS 59.8 52.8 35.5 19.8 13.4 56.1
SC 85.6 81.7 46.0 8.6 16.2 21.4
VCA-TB Base 84.7 79.5 17.5 29.9 6.1 7.4
CoT 85.9 85.3 3.0 20.3 3.1 8.2
MS 59.7 51.9 45.0 17.3 28.4 50.8
SC 85.3 79.9 48.0 11.3 25.0 20.6
SSR Base 84.7 76.1 24.5 35.0 32.3 49.3
CoT 84.9 82.3 10.0 26.0 22.8 54.5
MS 59.9 51.6 42.0 19.8 24.0 75.9
SC 85.1 78.7 64.0 10.1 29.7 61.1
Ty. Base 84.1 65.0 51.0 37.5 30.1 62.7
CoT 85.8 82.8 15.5 19.0 38.9 68.9
MS 59.0 50.0 48.5 18.6 31.0 81.0
SC 84.9 78.1 73.5 9.2 36.2 65.7
Avg. - 6.1 33.7 20.9 21.3 43.6
@@
Self-Probing (SP) is an alternative CEM (See prompt example in Appendix T for a description) that divides the process of generating an answer and confidence score into two steps. An initial prompt is used to have the model elicit only an answer (plus any corresponding rationale), and then for the second step, the model is prompted with this answer and is asked to provide its confidence based on the answer’s veracity without being informed that it originally generated the answer. By dividing the process of generating an answer and a confidence score in two different steps, it provides a potential buffer to the attack algorithm to reducing confidence scores when targeting a user query, possibly improving attack robustness. In Table 13, we test this assertion across Llama-3-8B and GPT-3.5. We find that across the different attack algorithms the performance of SP is comparable with the other elicitation methods, and has less robustness than the simple Base method for Llama-3, meaning in its basic form it does not serve as a strong defence method.
@@
We follow the approach of Kumar et al. [30] to study how closely aligned the verbalized numerical confidence scores are with the underlying model log probabilities for predicted tokens. Token-level probabilities provide the likelihood ${p(y_{i}|\bm{\mathbf{X}},\bm{\mathcal{P}})}$, of each individual word token $y_{i}$ in output sequence $\bm{\mathcal{Y}}$, can be aggregated (e.g., averaged) to derive an overall “confidence” which primarily represents the likelihood of the whole output (akin to language modelling loss), which is in contrast to verbal confidence which provides a singular number directly from output sequence $\bm{\mathcal{Y}}$. We focus on the Base CEM as it is the only one that produces a simple answer confidence score pair. The analysis is performed on Llama-3-8B as its internal states are accessible. We attain a confidence score using log probabilities by finding the original predicted multiple choice answer token (e.g., B) and taking its log probability. We find the maximal log probabilities for each of the remaining potential answer tokens capturing all potential variants (e.g., a, A) in the token space. We then run all maximal probabilities through a softmax function and use the highest value for the final log probability-based confidence score.
Table 15 shows the average confidence prior to and post attack and the corresponding average normalized confidence based on token log probabilities for each dataset and attack method. We also calculate the Spearman correlation coefficient similar to Kumar et al. [30] between the generated confidence scores and log probability-based scores both before and after undergoing adversarial attacks. The Spearman’s rank correlation coefficient measures the degree of association between two distributions of verbalized and log probability-based scores and as a result reveals how honest the model’s verbalized scores are with its underlying log probabilities. Despite the average verbalized and log probabilities confidence scores being close, we find that the correlation is weak or even negative in most cases and hence the log probabilities are not well aligned between the verbal confidence scores. The adversarial confidence scores in most cases show even poorer alignment than the original confidence scores despite both confidence measures decreasing compared to pre-attack, meaning additional misalignment is introduced through the confidence attacks.
@@
Verb. Cf. Verb. Adv. Cf Log Probs. Adv. Log Probs. Corr Adv. Corr

### v3

@@ C.1 ConfidenceTriggers Settings
To optimize the performance of ConfidenceTriggers, we evaluated a variety of algorithm parameters. However, given the significant amount of iterations necessary to optimize the algorithm, we limited some parameters like trigger prompt length while still aiming for strong results. Regarding our training text corpus, for MedMCQA and StrategyQA, we use 70% of the data for optimizing the triggers, and evaluate it on the remaining 30% (out of which 200 random samples are selected). Given the relatively smaller size of TruthfulQA, we train the triggers on questions with five multiple choices options (yielding 197 different examples) and evaluate on the 203 questions with four options. For our trigger prompt length $\beta$ we choose 20 tokens, and for the total number of prompts we attempt to optimize $\alpha$ we also choose 20. We experimented with optimizing between 20 and 50 total generations. For the number of samples for the initial estimation $S$, we choose 200, and the value $\xi$ for each iteration we choose 12, leading to 48 different sampled training cases used for estimation for each prompt each generation. We choose a fifth of prompts (i.e., 4) with the lowest loss to be elites. An example of the generated triggers can be found in Appendix X.
The InitializePrompts($\alpha,\beta$) function in the main algorithm that is used to create the initial set of $\alpha$ triggers before optimization occurs, essentially for each trigger generates a string of a $\beta$ number of random tokens that are uniformly sampled from the set of confidence related words. The AverageConfidence function is defined in Algorithm 5.
Our procedure for Cross-Over and Mutation functions remains largely faithful to lapid2023open and are standard implementations for these functions given two parent sequences of data. Cross-Over is the standard Single point Cross-Over [10.7551/mitpress/1090.001.0001] operation and Mutation is Random (uniform) Mutation [10.5555/534133] where words at randomly chosen indices are replaced. Pseudo code for Cross-Over and Mutation can be found in Algorithm 3 and 4 respectively. The tournament select process is top-K like in 10.1162/evco.1996.4.4.361.
@@
[…]
@@
Algorithm 5 ConfidenceTriggers AverageConfidence Function
1:Input: Confidence Estimation Method $z$, # Samples for Estimation $n$, Prompt Set $\mathbf{P}$
2:Optional: Trigger string $t$
3:Output: Average confidence $C$
[…]
5:if $t$ then
6:  $\mathbf{Confidences_{n}}\leftarrow$ Generate confidence using $z$ for $p\in\mathbf{P_{n}}$ with $t$ appended to $p$
7:else
8:  $\mathbf{Confidences_{n}}\leftarrow$ Generate confidence using $z$ for $p\in\mathbf{P_{n}}$
9:end if
[…]
@@ Appendix D ConfidenceTriggers-AutoDAN Details
In this section, we present the algorithm along with the details for the AutoDAN variation of the ConfidenceTriggers algorithm. The general algorithm can be found in Algorithm 6 and generally follows the original version of ConfidenceTriggers but is modified to incorporate the steps and methods of the AutoDAN-HGA algorithm in liu2024autodan, in that we make use of the hierarchal structure of prompts and incorporate both sentence and paragraph level optimizations.
@@
Momentum Word Score is calculated the same way as in AutoDAN-HGA. A dictionary is established in which words from each of the prompts are added if they are not part of a collection of common stop words, punctuation, proper nouns etc. When a word is added as a key its score is added as a value. The score is derived from the loss of the prompt it originally came from. If the same word is first being added and it is present in multiple prompts, the score is averaged. If a word has a key it is updated using a momentum mechanism, whereby the updated value is averaged between its current (previous) value and the newly computed $(\mbox{current value}+\mbox{new value})/2$. The words in the dictionary are finally ranked based on their scores in descending order.
@@
A total of 100 generations is used to optimize the prompts. The remaining settings are consistent with those detailed in Appendix C.1. An example of the final generated triggers can be found in Appendix X.
@@
5:  $\mathcal{L}_{z}$ $\leftarrow$ AverageConfidence($z$, $S$) $\triangleright$ Base confidence w/o any trigger
@@
7:$P$ $\leftarrow$ InitializePrompts($\alpha$)
@@
12:     $\mathcal{L}_{p_{g}}$ $\leftarrow\mathcal{L}_{p_{g}}\mathbin{\|}$ AverageConfidence($p$, $z$, $\xi$) - $\mathcal{L}_{z}$
@@
31:    $p_{a}$ $\leftarrow$ Mutation($p_{a}$)
32:    $p_{b}$ $\leftarrow$ Mutation($p_{b}$)
@@
35:  if $|\mathbf{P}_{g}|$ >> $|\mathbf{P}|$ then
@@
38:  $\mathbf{P}$ $\leftarrow$ $\mathbf{P}_{g}$
@@
[…]
@@ Appendix E Confidence Agnostic Attacks
SC-CA 94.6 90.7 100.0 85.5 64.1
We examine if CEMs are robust against generic confidence-agnostic adversarial attacks. To this end, we compare the confidences generated for the original instances in SST-2 [socher-etal-2013-recursive] and their AdvGLUE [wang2021adversarial] versions since these comprehensive adversarial modifications were solely optimized for misclassifications. The results are in Table 8, where the difference in confidences are minor, with at most a 5% drop despite a high percentage of samples being effected. CEMs are shown to be generally robust against confidence-agnostic perturbations, even though they are effective in reducing accuracy by 15-20%.
@@ Appendix F Full Results Tables
SC 91.2 88.5 47.5 5.8 20.5 67.6 91.2 88.4 40.0 6.9 16.1 37.5
SC-CA 91.6 87.1 64.0 7.1 18.0 64.1 91.1 87.9 44.5 6.1 11.7 37.9
SQA Base 88.6 82.5 51.0 12.0 47.6 61.8 88.6 82.9 49.0 11.6 29.9 41.7
CoT 91.5 83.8 51.0 15.0 44.6 55.9 91.6 85.9 38.5 14.9 16.4 28.6
MS 80.1 64.9 66.5 22.8 44.7 30.0 79.5 65.7 66.5 20.6 27.4 34.4
SC 90.1 80.0 65.5 15.4 41.3 42.4 89.7 81.5 55.5 14.8 32.5 35.3
SC-CA 91.9 84.4 62.5 12.0 37.0 61.5 91.6 86.0 58.5 9.7 20.5 42.6
Avg. - 5.7 48.0 11.4 27.4 53.6 - 4.8 39.4 10.9 16.3 31.2
Table 11: Results using SSR user query attack on TQA using: Llama-3.1-8B, Llama-3.3-70B, and o3-mini. We observe that the results are similar to the original models in our study, and the same conclusions can be drawn.
Cf. Adv. Cf. % Aff. #Iter. Δ Aff. Cf. %Flp(OC) % Flp(OW)
Llama-3.1-8B Base 81.5 76.8 27.0 7.3 17.2 14.3 42.2
CoT 87.2 83.3 32.5 7.1 11.9 30.6 54.9
MS 71.7 59.6 67.0 7.6 18.0 16.3 60.5
SC-CA 85.2 80.7 39.5 7.2 11.3 32.3 52.5
Llama-3.3-70B Base 92.2 88.8 19.5 9.1 17.1 25.0 39.5
CoT 89.8 85.8 20.0 8.8 19.6 24.2 51.3
MS 71.1 60.6 66.0 7.3 15.8 13.2 43.8
SC-CA 90.0 86.4 32.5 8.4 10.9 19.0 40.5
o3-mini Base 89.8 87.6 34.0 7.9 6.4 21.9 59.7
CoT 91.0 85.6 54.5 7.1 10.0 18.8 58.2
MS 72.5 63.1 75.5 7.0 12.4 19.9 59.4
SC-CA 89.5 86.0 53.5 7.5 5.3 9.4 49.0
@@ Appendix H Confidence Phrases
Table 12: Effectiveness of different phrases appended to user queries on TruthfulQA. Averages Δ Overall Cf. are shown for cases where confidence goes up and down respectively. Negative values represent increases in average confidence compared to the original examples. Averages for Δ Overall Cf. are separated by where the confidence goes down and up respectively.
@@
Phrase Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW) Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
@@
SC 57.0 31.2 97.5 32.0 21.8 20.2 54.3 39.7 99.5 39.9 4.6 44.9
SC-CA 64.8 21.8 97.5 22.4 69.7 76.6 54.0 39.4 99.5 39.6 3.1 44.4
Avg. - 21.0/-6.9 92.0 14.2 60.7 47.4 26.7/-5.5 88.1 18.4 8.6 31.6
Table 13: Effectiveness of different phrases appended to system prompts on TruthfulQA. Averages Δ Overall Cf. are shown for cases where confidence goes up and down respectively. Negative values represent increases in average confidence compared to the original examples. Averages for Δ Overall Cf. are separated by cases where the confidence goes down and up respectively.
@@
Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW) Adv. Cf. Δ Overall Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
@@
SC 81.2 7.0 86.0 8.1 6.7 20.9 88.5 5.2 83.0 6.2 5.6 21.3
SC-CA 82.2 4.8 91.5 5.3 86.8 74.7 88.9 4.5 82.5 5.4 4.7 16.9
Avg. - 8.2/-3.5 74.9 5.1 55.3 48.6 6.1/-1.5 65.1 5.4 6.4 21.7
Overall, we see that model behaviour when prompted through user queries varies greatly, but generally responds to the requests and results in great differences in confidence, particularly with the Base CEM. CoT and SC are in general overconfident compared to all methods regardless of how the model is prompted. The level of affected samples is very high, but is not 100% in the majority of cases, hence even this method of directly prompting the model is not always effective at changing model confidence behaviour. Targeting the system prompt is less effective than targeting the user query, particularly with GPT-3.5. The models are likely in most cases ignoring the prompt contents as it is not core to their task.
@@ Appendix I Confidence Expressed Via Likert Scale
Cf. Adv. Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B VCA-TF Base 77.9 73.9 7.5 53.3 1.9 9.8
CoT 87.1 84.4 6.0 45.0 10.5 13.2
MS 31.6 23.6 42.0 19.0 19.5 33.6
SC 81.8 74.8 34.5 20.2 21.2 18.8
SC-CA 84.8 77.8 34.5 20.5 39.4 26.7
VCA-TB Base 77.9 73.6 8.0 54.7 2.8 9.8
CoT 87.1 82.3 12.0 39.8 15.1 14.0
MS 31.6 22.1 48.5 19.6 24.1 35.4
SC 84.3 74.0 41.0 25.1 29.3 25.7
SC-CA 85.2 76.1 40.0 22.8 32.3 31.7
SSR Base 77.9 70.4 17.0 44.6 25.9 39.1
CoT 87.1 78.3 22.5 39.4 22.1 50.9
MS 31.6 21.7 56.0 17.6 29.9 54.0
SC 81.9 67.8 64.5 21.9 26.0 66.3
SC-CA 84.3 73.0 61.0 18.4 28.0 57.9
Ty. Base 77.9 74.9 7.0 43.8 5.6 21.7
CoT 87.1 84.5 5.5 47.9 10.5 20.2
MS 31.6 27.8 27.5 13.7 11.5 25.7
SC 82.8 67.5 72.0 21.3 38.6 74.7
SC-CA 84.0 80.7 22.0 15.3 14.9 26.4
Avg. - 7.3 31.5 30.2 20.4 32.8
GPT-3.5 VCA-TF Base 84.5 80.6 12.5 31.3 3.1 8.5
CoT 85.9 85.5 2.0 21.5 0.0 5.3
MS 59.8 52.8 35.5 19.8 13.4 56.1
SC 86.1 82.1 45.5 8.7 22.0 24.7
SC-CA 85.6 81.7 46.0 8.6 16.2 21.4
VCA-TB Base 84.7 79.5 17.5 29.9 6.1 7.4
CoT 85.9 85.3 3.0 20.3 3.1 8.2
MS 59.7 51.9 45.0 17.3 28.4 50.8
SC 86.2 81.2 50.5 9.8 23.6 20.5
SC-CA 85.3 79.9 48.0 11.3 25.0 20.6
SSR Base 84.7 76.1 24.5 35.0 32.3 49.3
CoT 84.9 82.3 10.0 26.0 22.8 54.5
MS 59.9 51.6 42.0 19.8 24.0 75.9
SC 86.5 80.5 58.0 10.3 31.5 64.3
SC-CA 85.1 78.7 64.0 10.1 29.7 61.1
Ty. Base 84.1 65.0 51.0 37.5 30.1 62.7
CoT 85.8 82.8 15.5 19.0 38.9 68.9
MS 59.0 50.0 48.5 18.6 31.0 81.0
SC 85.8 78.6 74.0 9.8 31.8 62.0
SC-CA 84.9 78.1 73.5 9.2 36.2 65.7
Avg. - 6.0 38.3 18.7 22.5 43.4
In addition to numerical confidence scores, it is possible to prompt the model to produce confidence scores in the form of phrases (e.g., fairly certain) derived from a Likert scale, which can potentially alter model behaviour and bias compared to numeric scores given that phrases are more naturalistic and resemble how humans typically express confidence in the real-world. We aim to analyze how attack performance changes when a model is asked to provide a confidence based on a seven point Likert scale where the options are the following:
@@
The statements are converted to numerical confidence scores for use in attack algorithms by interpolation. The first option is set to max confidence (100%) and the last to maximum entropy ($1/num\_mc\_options*100\%$), and the remaining values are interpolated between these two values according to a linear scale. We determine how effective each of our attack algorithms are in this context, and Table 14 presents the key results. Overall, there is not a noticeable improvement in attack performance compared to when the LLMs produce numerical scores, and performance is very similar in terms of average confidence drop and percent of samples that undergo a drop. We do see that Δ Aff. Cf. is very high but this can be attributed due to the discrete nature of the scale meaning any change in answer results in a large confidence drop (i.e., it cannot be 5% like when using numerical verbalized scores). In general, however, we do see a similar level of invariance in terms of confidence here and as such this behaviour of the model generalizes across both numerical and phrase-like confidence estimates which means LLMs model both numerical and phrasal confidence similarly and present rigidness in both.
@@
Self-Probing (SP) is an alternative CEM (See prompt example in Appendix W for a description) that divides the process of generating an answer and confidence score into two steps. An initial prompt is used to have the model elicit only an answer (with the addition of any corresponding rationale), and then for the second step, the model is prompted with this answer and is asked to provide its confidence based on the answer’s veracity without being informed that it originally generated the answer. By dividing the process of generating an answer and a confidence score in two different steps, it provides a potential buffer to the attack algorithm to reducing confidence scores when targeting a user query, possibly improving attack robustness. In Table 15, we test this assertion across Llama-3-8B and GPT-3.5. We find that across the different attack algorithms the performance of SP is comparable with the other elicitation methods, and has less robustness than the simple Base method for Llama-3, meaning in its basic form it does not serve as a strong defence method.
@@
Verb. Cf. Verb. Adv. Cf. Log Probs. Adv. Log Probs. Corr. Adv. Corr.
@@
We follow the approach of kumar-etal-2024-confidence to study how closely aligned the verbalized numerical confidence scores are with the underlying model log probabilities for predicted tokens. Token-level probabilities provide the likelihood $p(y_{i}|\bm{\mathbf{X}},\bm{\mathcal{P}})$}, of each individual word token $y_{i}$ in output sequence $\bm{\mathcal{Y}}$, can be aggregated (e.g., averaged) to derive an overall “confidence” which primarily represents the likelihood of the whole output (akin to language modelling loss), which is in contrast to verbal confidence which provides a singular number directly from output sequence $\bm{\mathcal{Y}}$. We focus on the Base CEM as it is the only one that produces a simple answer confidence score pair. The analysis is performed on Llama-3-8B as its internal states are accessible. We attain a confidence score using log probabilities by finding the original predicted multiple choice answer token (e.g., B) and taking its log probability. We find the maximal log probabilities for each of the remaining potential answer tokens capturing all potential variants (e.g., a, A) in the token space. We then run all maximal probabilities through a softmax function and use the highest value for the final log probability-based confidence score.
Table 17 shows the average confidence prior to and post attack and the corresponding average normalized confidence based on token log probabilities for each dataset and attack method. We also calculate the Spearman correlation coefficient similar to kumar-etal-2024-confidence between the generated confidence scores and log probability-based scores both before and after undergoing adversarial attacks. The Spearman’s rank correlation coefficient measures the degree of association between two distributions of verbalized and log probability-based scores and as a result reveals how honest the model’s verbalized scores are with its underlying log probabilities. Despite the average verbalized and log probabilities confidence scores being close, we find that the correlation is weak or even negative in most cases and hence the log probabilities are not well aligned between the verbal confidence scores. The adversarial confidence scores in most cases show even poorer alignment than the original confidence scores despite both confidence measures decreasing compared to pre-attack, meaning additional misalignment is introduced through the confidence attacks.